Find the volume of the solid generated by revolving the region bounded by the curves y = x^2 and y = 4x − x^2 about the line y = 6.
1
Expert's answer
2020-06-30T17:27:55-0400

Solution:

Solve a system

y=x2

y=4x-x2

We obtain x=0,x=2

Then in a new coordinates y'=y-6 we have two functions:y'1=x2-6, y'2=4x-x2-6

V="\pi" "\int" 20((y'1)2-(y'2)2)dx=...=

π"\int" 20(8x3-40x2+48x)dx="\pi" (2x4-40x3/3+24x2)|20="\pi" (32-320/3+96)=64"\pi"/3